Differentiate the following functions


  1. y=(x2+5x+2)3
  2. y=(2x2-x+3)5
  3. y=(x2-2x+6)7
  4. ∛4x-5
  5. ∛6-7x
1
Expert's answer
2022-04-13T16:28:07-0400
  1. "y'=3(x^2+5x+2)^2*(x^2+5x+2)'=3(x^2+5x+2)^2(2x+5)"
  2. "y'=5(2x^2-x+3)^4*(2x^2-x+3)'=5(2x^2-x+3)^4(4x-1)"
  3. "y'=7(x^2-2x+6)^6*(x^2-2x+6)'=7(x^2-2x+6)^6(2x-2)"
  4. "y'={\frac 1 {3*(4x-5)^{2/3}}}*(4x-5)'={\frac 4 {3(4x-5)^{2/3}}}"
  5. "y'={\frac 1 {3(6-7x)^{2/3}}}*(6-7x)'={\frac {-7} {3(6-7x)^{2/3}}}"